OpenStudy (anonymous):

a numbr consists of two digits and the digit in the ten's place exceeds that in the unit's place by 5. it 5 times the sum of the digits be subtracted from the number are reversed. then the sum of digits of the numbr is

OpenStudy (anonymous):

lets say the unit's place number is x then the 10s place digit is x+5 so the number is 10(x+5)+x. correct?

OpenStudy (anonymous):

explain me

OpenStudy (anonymous):

\[if \tan \theta=1 then the value of 8\sin \theta+5\cos \theta/\sin ^{3}\theta-2\cos ^{3}\theta+7\cos \theta is.............. explain me\]